Find the slope of the line that goes through the given points
(2,7) and (5,-8)

Answer by stanbon(75887)

Find the slope of the line that goes through the given points
(2,7) and (5,-8)
------
slope = (-8-7)/(5-2) =-15/3 = -5
==================
Cheers,
Stan H.
==================